A plate with positive charge is 0.05 m higher than a plate with negative charge, and the electric field has an intensity of E=5x10^4N/C. Determine the work performed by the electric field E when a charge of +5x10^-6C moves from the negative plate to the positive plate.
1
Expert's answer
2021-10-14T18:39:12-0400

By definition, the work required to move charge "q = 5\times 10^{-6}C" from A (negative plate) to B (positive plate) is given as follows:


"W = q\phi_{AB}"

where "\phi_{AB}" is the potential difference between A and B. In turn, the potential difference between capacitor plates is given as follows:


"\phi_{AB} = -Ed"

where "E=5\times 10^4N/C" is the electric field, and "d = 0.05m" is the distance between plates ('-' sign since the charge moves from negative to positive). Thus, obtain:


"W = -qEd = -5\times 10^{-6}\cdot 5\times 10^4\cdot 0.05 = -0.0125 N\cdot m"

Answer. -0.0125 Nm.